# Deflated Dynamics Value Iteration

### v1

@@ 5.1.  Implementation of DDTD
10:     $W(X_{k})\leftarrow W(X_{k})+\eta_{k}\big{[}\alpha R_{k}+\alpha\gamma V(X_{k}^{\prime})-\alpha\gamma(E_{s}V)(X_{k})+(1-\alpha)V(X_{k})-W(X_{k})\big{]}$
@@
Figure 1: Comparison of DDVI with different ranks, AutoPI, and AutoQR against VI in (left) Chain Walk and (right) Maze. Rate of DDVI with AutoPI and AutoQR changes when $E_{s}$ is updated.
@@
6.  Experiments
For our experiments, we use the following environments: Maze with $5\times 5$ states and $4$ actions, Cliffwalk with $3\times 7$ states with 4 actions, Chain Walk with 50 states with 2 actions, and random Garnet MDPs [Bhatnagar et al., 2009] with 200 states. The discount factor is set to $\gamma=0.99$ for DDTD experiments and $\gamma=0.995$ in other experiments. Appendix 10 provides full definitions of the environments and policies we used. All experiments were carried out on local CPUs. For comparisons, we use the normalized error of $V^{k}$ defined as $\|V^{k}-V^{\pi}\|_{1}/\|V^{\pi}\|_{1}$.
DDVI with AutoPI, AutoQR, different fixed ranks.
In Figure 1, we compare rank-$s$ DDVI for multiple values of $s$ and DDVI with AutoPI and AutoQR against the VI in Chain Walk and Maze. QR iteration (Section 2) is used to calculate $E_{s}$ with Schur vectors. In almost all cases, DDVI exhibits a significantly faster convergence rate compared to VI. Aligned with the theory, we observe that higher ranks achieve better convergence rates. Also, as DDVI with AutoPI and AutoQR progress and update $E_{s}$, their convergence rate improves.
DDVI and other baselines.
We perform an extensive comparison of DDVI against the prior accelerated VI methods: Safe Accelerated VI (S-AVI)[Goyal and Grand-Clément, 2022], Anderson VI [Geist and Scherrer, 2018], PID VI [Farahmand and Ghavamzadeh, 2021], and Anchored VI [Lee and Ryu, 2023]. In this experiment, we use the Implicitly Restarted Arnoldi Method [Lehoucq et al., 1998] from SciPy package to calculate the eigenvalues and eigenvectors for DDVI. Figure 2 (top-left) shows the convergence behaviour of the algorithms by iteration count in 20 randomly generated Garnet environments, where our algorithms outperform all the baselines. This comparison might not be fair as the amount of computation needed for each iteration of algorithms is not the same. Therefore, in Figure 2 (top-right), we compare the algorithms by wall clock time. It can be seen that rank-$2$ DDVI initially has to spend time to calculate $E_{s}$ before it starts to update the value function, but after the slow start, it has the fastest rate. The fast rate can compensate for the initial time if a high accuracy is needed. DDVI with AutoQR and rank-$1$ DDVI show a fast convergence from the beginning.
We further investigate how the algorithms scale with the size of MDP and the discount factor. Figure 2 (bottom-left) shows the runtime to reach a normalized error of $10^{-4}$ against the number of states. DDVI with AutoQR and rank-$1$ DDVI show the best scaling. Note that even rank-$2$ calculates a non-trivial $E_{s}$ also scales competitively. Figure 2 (bottom-right) show the scaling behaviour with horizon of the MDP, which is $1/(1-\gamma)$. We measure the runtime to reach a normalized error of $10^{-4}$ for the horizon ranging from $100$ to $1000$ which corresponds to $\gamma$ ranging from $0.99$ to $0.999$. Remarkably, DDVI algorithms have a low constant runtime even with long horizon tasks with $\gamma\approx 1$. This is aligned with our theoretical result, as the rate $\gamma|\lambda_{s+1}|$ remains small as $\gamma$ approaches $1$.
Rank-$s$ DDTD with QR iteration.
We compare DDTD with TD Learning and Dyna. For the sake of making the setting more realistic, we consider the case where the approximate model $\hat{\mathcal{P}}$ cannot exactly learn the true dynamics. Figure 3 shows that DDTD with large enough rank, can outperform TD. Note that unlike Dyna, DDTD does not suffer from model error, which shows that DDTD only uses the model for acceleration and is not a pure model-based algorithm.
7.  Conclusion
In this work, we propose a framework for accelerating VI through matrix deflation. We theoretically analyzed our proposed methods DDVI and DDTD, and presented experimental results showing speedups in various setups.
The positive experimental results demonstrate matrix deflation to be a promising technique that may be applicable to a broader range of RL algorithms. One direction of future work is to extend the theoretical analysis DDVI for Control using a general rank-$s$ deflation matrix. Theoretically analyzing other RL methods combined with matrix deflation is another interesting direction.
Acknowledgements
AMF acknowledges the funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant program (2021-03701) and the Canada CIFAR AI Chairs program.
Appendix
7.1.  Prior works
@@
Prioritized sweeping and its several variants [Moore and Atkeson, 1993, Peng and Williams, 1993, McMahan and Gordon, 2005, Wingate et al., 2005, Andre et al., 1997, Dai et al., 2011] specify the order of asynchronous value function updates, which may lead to accelerated convergence to the true value function. Speedy Q-learning [Azar et al., 2011] changes the update rule of Q-learning and employs aggressive learning rates to accelerate convergence. Sidford et al. [2023] accelerate the computation of $\mathcal{P}^{\pi}V^{k}$ by sampling it with variance reduction techniques. Recently, there has been a growing body of research that explores the application of acceleration techniques of other areas of applied math to planning and RL: Geist and Scherrer [2018], Sun et al. [2021], Ermis and Yang [2020], Park et al. [2022], Ermis et al. [2021], Shi et al. [2019] applies Anderson acceleration of fixed-point iterations, Lee and Ryu [2023] applies Anchor acceleration of minimax optimization, Vieillard et al. [2020], Goyal and Grand-Clément [2022], Grand-Clément [2021], Bowen et al. [2021], Akian et al. [2022] applies Nesterov acceleration of convex optimization, and Farahmand and Ghavamzadeh [2021] applies ideas inspired by PID controllers in control theory.
@@
Deflation techniques were first developed for eigenvalue computation [Meirovitch, 1980, Saad, 2011, Golub and Van Loan, 2013]. Matrix deflation, eliminating top eigenvalues of a given matrix with leaving the rest of the eigenvalues untouched, has been applied to various algorithms such as conjugate gradient algorithm [Saad et al., 2000], principal component analysis [Mackey, 2008], generalized minimal residual method [Morgan, 1995], and nonlinear fixed point iteration [Shroff and Keller, 1993] for improvement of convergence.
Some prior work in RL has explored subtracting constants from the iterates of VI, resulting in methods that resemble our rank-$1$ DDVI. For discounted MDP, Devraj and Meyn [2021] propose relative Q-learning, which can roughly be seen as the Q-learning version of rank-1 DDVI for Control, and Bertsekas [1995] proposes an extrapolation method for VI, which can be seen as rank-1 DDVI with shur deflation matrix for PE. White [1963] introduced relative value iteration in the context of undiscounted MDP with average reward, and several variants were proposed [Gupta et al., 2015, Sharma et al., 2020].
@@
Matrix splitting has been studied in the RL literature to obtain an acceleration. Hastings [1968] and Kushner and Kleinman [1968] first suggested Gausss-Seidel iteration for computing value function. Kushner and Kleinman [1971] and Reetz [1973] applied Successive Over-Relaxation to VI and generalized Jacobi and Gauss-Seidel iteration. Porteus [1975] proposed several transformations of MDP that can be seen as a Gauss–Seidel variant of VI. Rakhsha et al. [2022] used matrix splitting with the approximate dynamics of the environment and extended to nonlinear operator splitting. Bacon and Precup [2016], Bacon [2018] analyzed planning with options through the matrix splitting perspective.
@@
Jaakkola et al. [1993] first proved the convergence of TD learning using the stochastic approximation (SA) technique. Borkar and Meyn [2000], Borkar [1998] suggested ODE-based framework to provide asymptotic convergence of SA including TD learning with an asynchronous update. Lakshminarayanan and Szepesvari [2018] study linear SA under i.i.d noise with respect to mean square error and Chen et al. [2020] study asymptotic convergence rate of linear SA under Markovian noise. Finite time analysis of TD learning first provided by Dalal et al. [2018] and extended to Markovian noise setting by Bhandari et al. [2018]. Leveraging Lyapunov theory, Chen et al. [2023] establishes finite time analysis of Markovian SA with an explicit bound.
8.  Proof of Theoretical Results
8.1.  Proof of Theorem 3.1
By definition of DDVI (2) and fixed point $V^{\pi}$, we have
$V^{k}-V^{\pi}$
$=(I-\alpha\gamma E_{s})^{-1}((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E_{s}))(V^{k-1}-V^{\pi})$
$=(I-\alpha\gamma E_{s})^{-1}((1-\alpha)(I-\alpha\gamma E_{s})^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1})((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E_{s}))(V^{k-2}-V^{\pi})$
$=(I-\alpha\gamma E_{s})^{-1}((1-\alpha)(I-\alpha\gamma E_{s})^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1})^{k-1}((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E_{s}))(V^{0}-V^{\pi}),$
@@ 8.  Proof of Theoretical Results
$\|V^{k}-V^{\pi}\|\leq C\left\|((1-\alpha)(I-\alpha\gamma E)^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E)(I-\alpha\gamma E)^{-1})^{k-1}\right\|\|V^{0}-V^{\pi}\|$
@@
First, suppose $E_{s}$ satisfies Facts 1 or 2. Let $U_{s}=[u_{1},\dots,u_{s}]$ and $V_{s}=[v_{1},\dots,v_{s}]$. Define $D_{s,f(\lambda_{i})}=\text{diag}(f(\lambda_{1}),\dots,f(\lambda_{s}))$ for some function $f$. Then $U_{s},V_{s},D_{s,f(\lambda_{i})}$ are $n\times s$ matrices and $E_{s}=U_{s}D_{s,\lambda_{i}}V_{s}^{\top}$. By Jordan decomposition, $\mathcal{P}^{\pi}=UJU^{-1}$ where $J$ is Jordan matrix with $J_{ii}=\lambda_{i}$ for $i=1,\dots,n$. Let $J_{s}$ be $s\times s$ submatrix of $J$ satisfying $(J_{s})_{ij}=J_{ij}$ for $1\leq i,j\leq s$. Then, by condition of Theorem 3.1, $J_{s}=\text{diag}(\lambda_{1},\dots,\lambda_{s})=D_{s,\lambda_{i}}$ and $i$-th column of $U$ is $u_{i}$ for $1\leq i\leq s$. By simple calculation, we get $\mathcal{P}^{\pi}U_{s}=U_{s}J_{s}=U_{s}D_{s,\lambda_{i}}$ and $(I-\alpha\gamma E_{s})^{-1}=I+U_{s}D_{s,\frac{\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}$.
@@
$(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1}$ $=\mathcal{P}^{\pi}-E_{s}.$
Then,
$=(1-\alpha)\left(I+U_{s}D_{s,\frac{\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}\right)+\alpha\gamma\left(UJU^{-1}-U_{s}D_{s,\lambda_{i}}V^{\top}_{s}\right)$
$=(1-\alpha)I+\alpha\gamma UJU^{-1}+U_{s}D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}$
$=U\left((1-\alpha)I+\alpha\gamma J+e_{1:s}D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V_{s}^{\top}U\right)U^{-1},$
where $e_{i}\in\mathbb{R}^{n}$ is $i$-th unit vector and $e_{1:s}=[e_{1},\dots,e_{s}]$, and $(1-\alpha)I+\alpha\gamma J+e_{1:s}D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V_{s}^{\top}U$ is upper traingular matrix with diagonal entries $\frac{(1-\alpha)}{1-\alpha\gamma\lambda_{1}},\dots,\frac{(1-\alpha)}{1-\alpha\gamma\lambda_{s}},(1-\alpha)+\alpha\gamma\lambda_{s+1},\dots,(1-\alpha)+\alpha\gamma\lambda_{n}$.
Now suppose $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}u_{i}^{\top}$ satisfies Fact  3. Similarly, let $U_{s}=[u_{1},\dots,u_{s}]$. Then, $E_{s}=U_{s}D_{s,\lambda_{i}}U^{\top}_{s}$. By Schur decomposition, $\mathcal{P}^{\pi}=URU^{\top}$ where $R$ is an upper triangular matrix with $R_{ii}=\lambda_{i}$ for $i=1,\dots,n$, and $U$ is a unitary matrix. By simple calculation, we have $\mathcal{P}^{\pi}U_{s}=U_{s}R_{s}$ where $R_{s}$ is $s\times s$ submatrix of $R$ such that $(R_{s})_{ij}=R_{ij}$ for $1\leq i,j\leq s$ and $(I-\alpha\gamma E_{s})^{-1}=I+U_{s}D_{s,\frac{\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}U^{\top}_{s}$.

### v2

@@ 5.1 Implementation of DDTD
10:     $W(X_{k})\leftarrow W(X_{k})+\eta_{k}(X_{k})\big{[}\alpha R_{k}+\alpha\gamma V(X_{k}^{\prime})-\alpha\gamma(E_{s}V)(X_{k})+(1-\alpha)V(X_{k})-W(X_{k})\big{]}$
@@
We note that DDTD simultaneously uses samples to learn the model and to update the value function, making the DDTD framework more effective than model-free learning and more computationally efficient than a purely model-based approach.
6 Experiments
For our experiments, we use the following environments: Maze with $5\times 5$ states and $4$ actions, Cliffwalk with $3\times 7$ states with 4 actions, Chain Walk with 50 states with 2 actions, and random Garnet MDPs (Bhatnagar et al., 2009) with 200 states. The discount factor is set to $\gamma=0.99$ for the comparison of DDVI with different ranks and the DDTD experiments, and $\gamma=0.995$ in other experiments. Appendix D provides full definitions of the environments and policies used for PE. All experiments were carried out on local CPUs. We report the normalized error of $V^{k}$ defined as $\|V^{k}-V^{\pi}\|_{1}/\|V^{\pi}\|_{1}$. [^2] The source code for the experiments can be found at https://github.com/adaptive-agents-lab/ddvi.
DDVI with AutoPI, AutoQR, different fixed ranks.
In Figure 1, we compare rank-$s$ DDVI for multiple values of $s$ and DDVI with AutoPI and AutoQR against the VI in Chain Walk and Maze. In our plots, we incorporate the cost of the QR iteration for a fair comparison. We consider the cost of a QR iteration to be $msC$ where $m$ is the number of QR iterations we applied, $s$ is the rank of DDVI, and $C$ is the cost of VI per iteration, and plot rank-$s$ DDVI starting from $ms$ iterations in Figures 1. In this experiment, QR iteration (Section 2) is used $100$ times to calculate $E_{s}$ with Schur vectors. In almost all cases, DDVI exhibits a significantly faster convergence rate compared to VI. Aligned with the theory, we observe that higher ranks achieve better convergence rates. Also, as DDVI with AutoPI and AutoQR progress and update $E_{s}$, their convergence rate improves.
Figure 1: Comparison of DDVI with different ranks, AutoPI, and AutoQR against VI in (left) Maze and (right) Chain Walk. The plots for DDVI do not start at iteration 00 because the costs of computing $E_{s}$ through QR iterations are incorporated as rightward shifts. Rate of DDVI with AutoPI and AutoQR changes when $E_{s}$ is updated.
@@ 6 Experiments
DDVI and other baselines.
We perform an extensive comparison of DDVI against the prior accelerated VI methods: Safe Accelerated VI (S-AVI)(Goyal & Grand-Clément, 2022), Anderson VI (Geist & Scherrer, 2018), PID VI (Farahmand & Ghavamzadeh, 2021), and Anchored VI (Lee & Ryu, 2023). In this experiment, we use the Implicitly Restarted Arnoldi Method (Lehoucq et al., 1998) from SciPy package to calculate the eigenvalues and eigenvectors for DDVI. Figure 2 (top-left) shows the convergence behaviour of the algorithms by iteration count in 20 randomly generated Garnet environments, where our algorithms outperform all the baselines. This comparison might not be fair as the amount of computation needed for each iteration of algorithms is not the same. Therefore, in Figure 2 (top-right), we compare the algorithms by wall clock time. It can be seen that rank-$2$ DDVI initially has to spend time to calculate $E_{s}$ before it starts to update the value function, but after the slow start, it has the fastest rate. The fast rate can compensate for the initial time if a high accuracy is needed. DDVI with AutoQR and rank-$1$ DDVI show a fast convergence from the beginning.
We further investigate how the algorithms scale with the size of MDP and the discount factor. Figure 2 (bottom-left) shows the runtime to reach a normalized error of $10^{-4}$ as a function of the number of states. DDVI with AutoQR and rank-$1$ DDVI show the best scaling. Note that even rank-$2$, for which the calculation of $E_{s}$ is non-trivial, also scales competitively. Figure 2 (bottom-right) show the scaling behaviour with horizon of the MDP, which is $1/(1-\gamma)$. We measure the runtime to reach a normalized error of $10^{-4}$ for the horizon ranging from $100$ to $1000$ which corresponds to $\gamma$ ranging from $0.99$ to $0.999$. Remarkably, DDVI algorithms have a low constant runtime even with long horizon tasks with $\gamma\approx 1$. This is aligned with our theoretical result, as the rate $\gamma|\lambda_{s+1}|$ remains small as $\gamma$ approaches $1$.
Rank-$s$ DDTD with QR iteration.
We compare DDTD with TD Learning and Dyna (Sutton, 1990; Peng & Williams, 1993) which simultaneously use samples to learn the model and to update the value function. For the sake of making the setting more realistic, we consider the case where the approximate model $\hat{\mathcal{P}}$ cannot exactly learn the true dynamics. Figure 3 shows that DDTD with large enough rank can outperform TD. Note that unlike Dyna, DDTD does not suffer from model error, which shows that DDTD only uses the model for acceleration and is not a pure model-based algorithm.
@@
7 Conclusion
In this work, we proposed a framework for accelerating VI through matrix deflation and matrix splitting. We theoretically analyzed the proposed methods, DDVI and DDTD, and presented experimental results showing speedups in various setups. The positive experimental results demonstrate that matrix deflation to be a promising technique that may be applicable to a broader range of RL algorithms.
One direction of future work is to extend the theoretical analysis DDVI for Control using a general rank-$s$ deflation matrix. Theoretically analyzing other RL methods combined with matrix deflation is another interesting direction. Another direction is incorporating the DDVI framework with a function approximator, such as a deep neural network, to handle large-scale or continuous state and action spaces.
Acknowledgments
We thank the anonymous reviewers, the action editor, as well as the members of the Adaptive Agents (Adage) Lab, in particular Tyler Kastner, for their constructive feedback. JL and EKR were supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (No.RS-2024-00421203). AMF acknowledges the funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant program (2021-03701). Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
Appendix A Prior Works
@@
Prioritized sweeping and its several variants (Moore & Atkeson, 1993; Peng & Williams, 1993; McMahan & Gordon, 2005; Wingate et al., 2005; Andre et al., 1997; Dai et al., 2011) specify the order of asynchronous value function updates, which may lead to accelerated convergence to the true value function. Speedy Q-learning (Azar et al., 2011) changes the update rule of Q-learning and employs aggressive learning rates to accelerate the convergence. Sidford et al. (2023) accelerate the computation of $\mathcal{P}^{\pi}V^{k}$ by sampling it with variance reduction techniques.
Recently, there has been a growing body of research that explores the application of acceleration techniques of other areas of applied math to planning and RL: Geist & Scherrer (2018); Sun et al. (2021); Ermis & Yang (2020); Park et al. (2022); Ermis et al. (2021); Shi et al. (2019) apply Anderson acceleration of fixed-point iterations, Lee & Ryu (2023) apply Anchor acceleration of minimax optimization, Vieillard et al. (2020); Goyal & Grand-Clément (2022); Grand-Clément (2021); Bowen et al. (2021); Akian et al. (2022) apply Nesterov acceleration of convex optimization, and Farahmand & Ghavamzadeh (2021) apply ideas inspired by PID controllers in control theory.
Specifically, regarding the variants of VI used in our experiments, S-AVI (Goyal & Grand-Clément, 2022) exhibits $O(\gamma_{s}^{k})$-rate where $\gamma_{s}$ satisfies $\gamma\leq\gamma_{s}\leq 1$, PID VI (Farahmand & Ghavamzadeh, 2021) exhibits a guaranteed $O\left(\left(\frac{\sqrt{1+\gamma}-\sqrt{1-\gamma}}{\sqrt{1+\gamma}+\sqrt{1-\gamma}}\right)^{k}\right)$ rate under reversible MDP (and possibly faster, depending on the MDP), and Anchored VI (Lee & Ryu, 2023) exhibits $O\left(\frac{(\gamma^{-1}-\gamma)(1+2\gamma-\gamma^{k+1})}{(\gamma^{k+1})^{-1}-\gamma^{k+1}}\right)$-rate in terms of Bellman error.
@@
Deflation techniques were first developed for eigenvalue computation (Meirovitch, 1980; Saad, 2011; Golub & Van Loan, 2013). Matrix deflation, eliminating top eigenvalues of a given matrix with leaving the rest of the eigenvalues untouched, has been applied to various algorithms such as the conjugate gradient algorithm (Saad et al., 2000), principal component analysis (Mackey, 2008), the generalized minimal residual method (Morgan, 1995), and nonlinear fixed-point iteration (Shroff & Keller, 1993) to improve the convergence.
Some prior work in RL has explored subtracting constants from the iterates of VI, resulting in methods that resemble our rank-$1$ DDVI. For discounted MDP, Devraj & Meyn (2021) propose relative Q-learning, which can roughly be seen as the Q-learning version of rank-1 DDVI for Control, and Bertsekas (1995) proposes an extrapolation method for VI, which can be seen as rank-1 DDVI with Schur deflation matrix for PE. White (1963) introduced relative value iteration in the context of undiscounted MDP with average reward, and several variants were proposed (Gupta et al., 2015; Sharma et al., 2020).
@@
Matrix splitting has been studied in the RL literature to obtain an acceleration. Hastings (1968) and Kushner & Kleinman (1968) first suggested Gausss-Seidel iteration for computing value function. Kushner & Kleinman (1971) and Reetz (1973) applied Successive Over-Relaxation to VI and generalized Jacobi and Gauss-Seidel iterations. Porteus (1975) proposed several transformations of MDP that can be seen as a Gauss–Seidel variant of VI. Rakhsha et al. (2022) used matrix splitting with the approximate dynamics. Bacon & Precup (2016); Bacon (2018) analyzed planning with options through the matrix splitting perspective.
@@
Jaakkola et al. (1993) first proved the convergence of TD learning using the stochastic approximation (SA) technique. Borkar & Meyn (2000); Borkar (1998) suggested ODE-based framework to provide asymptotic convergence of SA including TD learning with an asynchronous update. Lakshminarayanan & Szepesvari (2018) study linear SA under i.i.d. noise with respect to mean square error, and Chen et al. (2020) study asymptotic convergence rate of linear SA under Markovian noise. Finite-time analysis of TD learning was first provided by Dalal et al. (2018) and extended to Markovian noise setting by Bhandari et al. (2018). Leveraging Lyapunov theory, Chen et al. (2023) establishes finite-time analysis of Markovian SA with an explicit bound.
Appendix B Proof of Theoretical Results
We prove the theoretical results in this appendix.
B.1 Proof of Theorem 3.1
By the definition of DDVI (equation 4) and the property of fixed point $V^{\pi}$, which satisfies $V^{\pi}=(I-\alpha\gamma E)^{-1}(\alpha r^{\pi}+((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E))V^{\pi})$ as mentioned in Section 3.2, we have
$V^{k}-V^{\pi}=$
$(I-\alpha\gamma E_{s})^{-1}\left[(1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})\right](V^{k-1}-V^{\pi})=$
$(I-\alpha\gamma E_{s})^{-1}\left[(1-\alpha)(I-\alpha\gamma E_{s})^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1}\right]((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E_{s}))(V^{k-2}-V^{\pi})=$
$(I-\alpha\gamma E_{s})^{-1}\left[(1-\alpha)(I-\alpha\gamma E_{s})^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1}\right]^{k-1}((1-\alpha)I+\alpha\gamma(\mathcal{P}^{\pi}-E_{s}))(V^{0}-V^{\pi}),$
@@ Appendix B Proof of Theoretical Results
$\|V^{k}-V^{\pi}\|\leq C\left\|((1-\alpha)(I-\alpha\gamma E)^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E)(I-\alpha\gamma E)^{-1})^{k-1}\right\|\|V^{0}-V^{\pi}\|,$ (9)
@@
First, suppose $E_{s}$ satisfies Facts 1 or 2. Let $U_{s}=[u_{1},\dots,u_{s}]$ and $V_{s}=[v_{1},\dots,v_{s}]$. Define $D_{s,f(\lambda_{i})}=\text{diag}(f(\lambda_{1}),\dots,f(\lambda_{s}))$ for some function $f$. Then $U_{s},V_{s},D_{s,f(\lambda_{i})}$ are $s\times s$ matrices and $E_{s}=U_{s}D_{s,\lambda_{i}}V_{s}^{\top}$. By Jordan decomposition, $\mathcal{P}^{\pi}=UJU^{-1}$ where $J$ is Jordan matrix with $J_{ii}=\lambda_{i}$ for $i=1,\dots,n$. Let $J_{s}$ be $s\times s$ submatrix of $J$ satisfying $(J_{s})_{ij}=J_{ij}$ for $1\leq i,j\leq s$. Then, by condition of Theorem 3.1, $J_{s}=\text{diag}(\lambda_{1},\dots,\lambda_{s})=D_{s,\lambda_{i}}$, and the $i$-th column of $U$ is $u_{i}$ for $1\leq i\leq s$. By simple calculation, we get $\mathcal{P}^{\pi}U_{s}=U_{s}J_{s}=U_{s}D_{s,\lambda_{i}}$ and $(I-\alpha\gamma E_{s})^{-1}=I+U_{s}D_{s,\frac{\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}$.
@@
$(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1}$ $=\mathcal{P}^{\pi}-E_{s}.$ (10)
Then, the term on the right-hand side (RHS) of equation 9 can be written as
$(1-\alpha)(I-\alpha\gamma E_{s})^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1}=$
$(1-\alpha)\left(I+U_{s}D_{s,\frac{\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}\right)+\alpha\gamma\left(UJU^{-1}-U_{s}D_{s,\lambda_{i}}V^{\top}_{s}\right)=$
$(1-\alpha)I+\alpha\gamma UJU^{-1}+U_{s}D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}=$
$U\left((1-\alpha)I+\alpha\gamma J+e_{1:s}D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V_{s}^{\top}U\right)U^{-1},$
where $e_{i}\in\mathbb{R}^{n}$ is the $i$-th unit vector and $e_{1:s}=[e_{1},\dots,e_{s}]$, and $(1-\alpha)I+\alpha\gamma J+e_{1:s}D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V_{s}^{\top}U$ is an upper triangular matrix with diagonal entries $\frac{(1-\alpha)}{1-\alpha\gamma\lambda_{1}},\dots,\frac{(1-\alpha)}{1-\alpha\gamma\lambda_{s}},(1-\alpha)+\alpha\gamma\lambda_{s+1},\dots,(1-\alpha)+\alpha\gamma\lambda_{n}$.
Now suppose $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}u_{i}^{\top}$ satisfies Fact 3. Similarly, let $U_{s}=[u_{1},\dots,u_{s}]$. Then, $E_{s}=U_{s}D_{s,\lambda_{i}}U^{\top}_{s}$. By Schur decomposition, $\mathcal{P}^{\pi}=URU^{\top}$ where $R$ is an upper triangular matrix with $R_{ii}=\lambda_{i}$ for $i=1,\dots,n$, and $U$ is a unitary matrix. By simple calculation, we have $\mathcal{P}^{\pi}U_{s}=U_{s}R_{s}$ where $R_{s}$ is the $s\times s$ submatrix of $R$ such that $(R_{s})_{ij}=R_{ij}$ for $1\leq i,j\leq s$ and $(I-\alpha\gamma E_{s})^{-1}=I+U_{s}D_{s,\frac{\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}U^{\top}_{s}$.